OpenStudy (anonymous):

use scientific notation to multiply the following two numbers. Express the number using scientific notation; retain all nonzero decimal places. (8.1 * 10^7) * (4.2 * 10^9)=

OpenStudy (unklerhaukus):

\[(8.1 \times 10^7) \times (4.2 \times 10^9)\]\[=(8.1 \times 4.2) \times (10^7 \times 10^9)\]\[=(8.1 \times 4.2)\times10^{7+9}\]\[=\]\[=\] you final answer should only one number before the decimal place

OpenStudy (anonymous):

my final answer should be one number before the decimal place?

OpenStudy (anonymous):

34.02* 10^16

OpenStudy (unklerhaukus):

well what i mean to say is \[23.4\times10^6\] is not in scientific notation because it should be \[2.34\times10^7\]

OpenStudy (anonymous):

ok, so my answer would be......3.402 * 10 ^16?

OpenStudy (unklerhaukus):

if you are moving the decimal place to the left you are effectively dividing by ten, to balance this you have to multiply by ten, ie add one to the exponent of ten

OpenStudy (anonymous):

ok, so that is who your exponent changed from 10^6 to 10^7 ?

OpenStudy (anonymous):

*why* not who

OpenStudy (unklerhaukus):

yeah, exactly \[34.02\times 10^{16}\longrightarrow3.402\times10^{16+1}\]

OpenStudy (anonymous):

thank you! that makes a lot more sense ;)